# Forward Ray Tracing and Hot Spots in Kerr Spacetime

### v1

@@ III.1.2 Light Echoes
which differs from the definition in Ref. Gralla and Lupsasca (2020a) by a factor of $2$. The non-integer number $n$ quantifies the length of the journey in the $\theta$ direction experienced by the photon. Every time $n$ increases by two, the photon experiences one more full oscillation in $\theta$. This definition can be connected to the integer definition of the number of half-orbits, $N$, through
@@ III.1.3 An Example
Figure 2: Contour plots illustrating the root-finding process in forward ray tracing for the example discussed in Sec. III.1.3: a point-like source located at $(r_{s}=10\,M$, $\theta_{s}=$$$ 90  °, $\phi_{s}=-$$)$ 45  ° ) outside a BH with spin $a=0.8\,M$. The observer is positioned at $(r_{o}=1000\,M$, $\theta_{o}=$$$ 17  °, $\phi_{o}=0)$. The yellow dashed lines at $\lambda=0$ divide the plane into regions with opposite azimuthal directions: $\phi_{f}-\phi_{s}>0$ (left) and $\phi_{f}-\phi_{s}<0$ (right). Black and red lines represent the solutions $\theta_{f}=\theta_{o}$ and $\phi_{f}=\phi_{o}+2k\pi$ ($k\in\mathbb{Z}$), respectively, and the blue points at their intersections correspond to the light ray solutions. Three sets of solutions are shown: direct emission (left), and lensed photons with contrasting signs of $\nu_{\theta}^{s}$ (middle and right).
@@
The fact that lensed images ($N\geq 1$) can be categorized into two sets is well expected and can be explained as follows. Lensed images are all formed by light rays carrying conserved quantities close to the critical curve. Their orbits, when propagating near the bound photon orbit, are nearly spherical. In the Schwarzschild case, nearly critical photons enter these quasi-bound orbits in two distinct directions: clockwise or counterclockwise. Therefore, we expect the initial photon directions of lensed images to naturally separate into two clusters. We find this phenomenon still widely holds in Kerr spacetime, despite the effect of spin that complicates the motion of photons. In the example here, the two initial photon clusters are distinguished by contrasting $\nu_{\theta}^{s}$, with odd and even level numbers $N$ in each category. Note that at the $N=7$ level, there are three different images, and therefore we label them as $7a$, $7b$, and $7c$. In fact, the appearance of multiple images with the same $N$ is common in Kerr spacetime, this will discussed further in Sec. III.3.
III.2 Time Delay and Position Angle Shift of Light Echoes
@@
where $\tau\approx 15\,M$ is the critical parameter characterizing the temporal period of a half-libration in the $\theta$ direction
Figure 3: Geodesics and images of the solutions found in forward ray tracing for the example discussed in Sec. III.1.3. Top left: Illustration of the image plane $\alpha-\beta$ showing the solutions of the lowest three levels. The dashed black line perpendicular to the plane represents the line of sight from the observer to the BH. The short black arrow on the BH denotes its spin direction, with its projection aligning with the $\beta$-axis on the plane. Light rays connect the point-like source (red sphere) with the faraway observer ($O$). Top right: Images from the point-like source on the image plane centered at the line-of-sight direction of the BH (black point). The direct image is shown in red, while the lensed images are in green for $\nu_{\theta}^{s}=+1$ and orange for $\nu_{\theta}^{s}=-1$. The critical curve is plotted in blue. Bottom: Trajectories of the light rays that form the images for $N$ from 00 to $9$.
@@ III.3 Image Level Degeneracy
As illustrated in the previous subsection, there can be multiple images at the same level $N$ from a point source. This phenomenon does not occur for a Schwarzschild BH, where geodesics at each level lie entirely within the same plane containing the point-like source, the BH, and the observer, sequentially projecting onto the image plane along a single line. In contrast, such cases become more common with increasing spin and inclination angle. This subsection delves into this phenomenon, termed image level degeneracy.
@@ IV Image Distortion in Kerr Spacetime
The forward ray tracing method developed in Sec. III enables the calculation of geodesics that connect a specific point near the BH to a distant observer. In reality, emission sources, such as hotspots, have a finite volume. This section addresses this practical aspect by introducing a perturbative mapping technique. This method relates small deviations near the original emission point to corresponding changes on the image plane. We formulate this perturbative mapping using conserved quantities to analyze the amplification both along and perpendicular to the critical curve for higher-order images. We then illustrate the scaling relations with an example in Schwarzschild spacetime, using straightforward geometric relations to calculate the amplification factor as a function of source position and image level.
@@ IV.2 Image Distortion in Schwarzschild Spacetime
Figure 9: Illustration of geometric relations for a spherical emission source ($S$) and its $N=1$ level image ($I$) on the image plane. $\rho_{I}$ and $\rho_{S}$ denote the distances from the BH projection on the image plane to $I$ and the projection of $S$ ($P_{S}$), respectively. The source center’s orbital plane aligns with the BH equatorial plane ($\theta=\pi/2$), and is located at $(r_{s},\phi_{s})$. The initial direction of geodesics from the source center is shown by the black arrow, forming an angle $\xi$ with the direction towards the BH. The two image edges perpendicular to the critical curve (CC) are marked by red dots, corresponding to points $A_{1}$ and $A_{2}$ on the spherical source.
To clearly illustrate the derivation of the amplification rate, we present the geometric relations for a radiating spherical hotspot in Fig. 9. The hotspot, denoted by $S$, has its center at $(r_{s},\,\theta_{s}=0,\,\phi_{s})$, with the initial direction of geodesics from this point indicated by the black arrow, leading to an image $I$ on the image plane. Although the example shown uses geodesics for $N=1$, the explanation applies to all image levels. The direct projection of the black hole on the image plane serves as the origin of the coordinate system, marked by a black dot, and the projection of the hotspot’s center is labeled as $P_{S}$. $\rho_{I}=\sqrt{\alpha^{2}+\beta^{2}}$ and $\rho_{S}$ represent the distances from the origin to $I$ and $P_{S}$, respectively. The gray contour line labeled ‘CC’ on the image plane denotes the critical curve.
@@ V Spacetime Tomography and Hotspot Localization
The forward ray tracing method is adept at simulating images of hotspots. This section demonstrates how this method can be applied to derive observational data, specifically for measuring the BH mass, spin, inclination angle, and the location of hotspots from observations of both the direct ($N=0$) and the first lensed ($N=1$) images of a hotspot. Since the intensity of higher-order images is exponentially suppressed, our focus will primarily be on these two initial images.
Emissions near a BH can exhibit significant time variability, which may be reflected across various image levels. The intensity of these images incorporates this variability, albeit with specific time delays. Thus, by correlating intensity fluctuations across different regions on the image plane with corresponding time delays, we can confirm their common origin as emissions from the same source Hadar et al. (2021). The time delay between the two lowest image levels from a point-like emission is sensitive to the BH mass, while their positional angle shift along the photon ring is sensitive to the BH spin Gralla and Lupsasca (2020a); Hadar et al. (2021); Andrianov et al. (2022). Additionally, the observer’s inclination angle to the BH and the hotspot’s spatial location can influence both the time delay and position angle shift, complicating the analysis as demonstrated in Sec. III.2. Forward ray tracing can mitigate these complexities by facilitating the scan over various source locations.
We model a point-like source located at $(r_{s},\theta_{s},\phi_{s})$ outside a Kerr BH with mass $M$ and spin $a$, directing sequential geodesics towards an observer positioned at an inclination angle $\theta_{o}$ relative to the BH. Without loss of generality, we set the observer’s azimuthal angle $\phi_{o}=0$. By successfully detecting both the direct emission ($N=0$) and the first lensed image ($N=1$) of a hotspot through correlated time variability of intensity, we can ascertain precise locations of these images on the image plane, represented by
@@
where $j=a,b,c,\ldots$ labels the $j$-th hotspot observed. Each selection of $\bm{\Theta}$ generates pairs of direct and the first lensed images $\bm{X}^{(j)}(\bm{\Theta})$ through the forward ray tracing method. These are then compared with the observed values $\mkern 1.0mu\overline{\mkern-1.0mu\bm{X}^{(j)}\mkern-1.0mu}\mkern 1.0mu$, which in the following two examples are assumed to match the theoretical values computed from the true parameters. Consequently, we introduce the likelihood function
@@
In scenarios where some variables in $\bm{\Theta}$ are pre-determined by other measurements, such as the inclination angle $\theta_{o}$, the spin projection PA, and the SMBH mass $M$, we employ the prior probability function
@@
where $l$ labels the pre-determined parameters with true values $\mkern 1.0mu\overline{\mkern-1.0mu\bm{\Theta}_{l}\mkern-1.0mu}\mkern 1.0mu$ and uncertainties $\sigma_{\bm{\Theta}_{l}}$. According to Bayes’ theorem, the posterior probability is then $p\left(\bm{\Theta}\middle|\bm{X}\right)\propto p\left(\bm{X}\middle|\bm{\Theta}\right)p(\bm{\Theta})$.
@@
V.1 M87 with Pre-determined PA and $\theta_{o}$
@@ VI Discussion
A direct application of this methodology is to BH imaging, particularly in connecting a point-like emission source outside a BH to a distant observer. This novel forward ray tracing approach proves more efficient than traditional backward ray tracing, especially in modeling hotspots and calculating higher-order images. By applying perturbative adjustments to the original geodesics around the emission point, our method can map a finite-size emission to distinct regions on the observer plane. Higher-order images, typically positioned near the critical curve, sequentially compress in the direction perpendicular to the curve while maintaining a stable amplification rate parallel to it. Consequently, the total area magnification rate exhibits exponential decay as the image level $N$ increases.
Our method is instrumental in spacetime tomography, particularly for measuring BH spin and mass. Utilizing observations from the EHT and ngEHT that capture both direct and lensed images, our approach can precisely determine BH characteristics. For M87, with predetermined spin projection and inclination angle, a single measurement suffices to resolve the BH’s mass, spin, and the emission location. In the case of Sgr A, two independent measurements of hotspot correlations are enough to determine all relevant parameters. Given the daily occurrence of flares from hotspots around Sgr A, the prospects for successful detection are promising, especially with ngEHT, which will offer tracking of hotspot motion Emami et al. (2023) and multi-frequency measurement capabilities Chael et al. (2023). The accuracy in determining universal spacetime parameters will improve as more data accumulates.
@@ Acknowledgements.
We are grateful to Xian Chen, Daniel D’Orazio, Shahar Hadar, Yosuke Mizuno, Daniel Palumbo, Diogo Ribeiro, Paul Tiede, Luka Vujeva, Bo Wang, George Wong, and Xiao Xue for useful discussions. L.Z. acknowledges financial support from Peking University for his visit to the Niels Bohr Institute. Z.Z. acknowledges financial support from China Scholarship Council (No. 202106040037). V.C. is a Villum Investigator and a DNRF Chair. We acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. We acknowledge financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855 and No 101131233.

### v2

@@ III.1.2 Light Echoes
which differs from the definition in Ref. Gralla and Lupsasca (2020a) by a factor of $2$. The noninteger number $n$ quantifies the length of the journey in the $\theta$ direction experienced by the photon. Every time $n$ increases by two, the photon experiences one more full oscillation in $\theta$. This definition can be connected to the integer definition of the number of half-orbits, $N$, through
@@ III.1.3 An Example
Figure 2: Contour plots illustrating the root-finding process in forward ray tracing for the example discussed in Sec. III.1.3: a pointlike source located at $(r_{s}=10\,M$, $\theta_{s}=$$$ 90  °, $\phi_{s}=-$$)$ 45  ° ) outside a BH with spin $a=0.8\,M$. The observer is positioned at $(r_{o}=1000\,M$, $\theta_{o}=$$$ 17  °, $\phi_{o}=0)$. The yellow dashed lines at $\lambda=0$ divide the plane into regions with opposite azimuthal directions: $\phi_{f}-\phi_{s}>0$ (left) and $\phi_{f}-\phi_{s}<0$ (right). Black and red lines represent the solutions $\theta_{f}=\theta_{o}$ and $\phi_{f}=\phi_{o}+2k\pi$ ($k\in\mathbb{Z}$), respectively, and the blue points at their intersections correspond to the light ray solutions. Three sets of solutions are shown: direct emission (left), and lensed photons with contrasting signs of $\nu_{\theta}^{s}$ (middle and right).
@@
Figure 3: Geodesics and images of the solutions found in forward ray tracing for the example discussed in Sec. III.1.3. Top left: Illustration of the image plane $\alpha-\beta$ showing the solutions of the lowest three levels. The dashed black line perpendicular to the plane represents the line of sight from the observer to the BH. The short black arrow on the BH denotes its spin direction, with its projection aligning with the $\beta$-axis on the plane. Light rays connect the pointlike source (red sphere) with the faraway observer ($O$). Top right: Images from the pointlike source on the image plane centered at the line-of-sight direction of the BH (black point). The direct image is shown in red, while the lensed images are in green for $\nu_{\theta}^{s}=+1$ and orange for $\nu_{\theta}^{s}=-1$. The critical curve is plotted in blue. Bottom: Trajectories of the light rays that form the images for $N$ from 00 to $9$.
The fact that lensed images ($N\geq 1$) can be categorized into two sets is well expected and can be explained as follows. Lensed images are all formed by light rays carrying conserved quantities close to the critical curve. Their orbits, when propagating near the bound photon orbit, are nearly spherical. In the Schwarzschild case, nearly critical photons enter these quasibound orbits in two distinct directions: clockwise or counterclockwise. Therefore, we expect the initial photon directions of lensed images to naturally separate into two clusters. We find this phenomenon still widely holds in Kerr spacetime, despite the effect of spin that complicates the motion of photons. In the example here, the two initial photon clusters are distinguished by contrasting $\nu_{\theta}^{s}$, with odd and even level numbers $N$ in each category. Note that at the $N=7$ level, there are three different images, and therefore we label them as $7a$, $7b$, and $7c$. In fact, the appearance of multiple images with the same $N$ is common in Kerr spacetime, which will discussed further in Sec. III.3.
III.2 Time Delay and Position Angle Shift of Light Echoes
@@ III.3 Image Level Degeneracy
As illustrated in the previous subsection, there can be multiple images at the same level $N$ from a point source. This phenomenon does not occur for a Schwarzschild BH, where geodesics at each level lie entirely within the same plane containing the pointlike source, the BH, and the observer, sequentially projecting onto the image plane along a single line. In contrast, such cases become more common with increasing spin and inclination angle. This subsection delves into this phenomenon, termed image level degeneracy.
@@ IV Image Distortion in Kerr Spacetime
The forward ray tracing method developed in Sec. III enables the calculation of geodesics that connect a specific point near the BH to a distant observer. In reality, emission sources, such as hotspots, have a finite volume. This section addresses this practical aspect by introducing a perturbative mapping technique Cunningham and Bardeen (1973); Ohanian (1987). This method relates small deviations near the original emission point to corresponding changes on the image plane. We formulate this perturbative mapping using conserved quantities to analyze the amplification both along and perpendicular to the critical curve for higher-order images. We then illustrate the scaling relations with an example in Schwarzschild spacetime, using straightforward geometric relations to calculate the amplification factor as a function of source position and image level.
@@ IV.2 Image Distortion in Schwarzschild Spacetime
Figure 9: Illustration of geometric relations for a spherical emission source ($S$) and its $N=1$ level image ($I$) on the image plane. $\rho_{I}$ and $\rho_{S}$ denote the distances from the BH projection on the image plane to $I$ and the projection of $S$ ($P_{S}$), respectively. The source center’s orbital plane aligns with the BH equatorial plane ($\theta=\pi/2$), and is located at $(r_{s},\phi_{s})$. The initial direction of geodesics from the source center is shown by the black arrow, forming an angle $\xi$ with the direction toward the BH. The two image edges perpendicular to the critical curve (CC) are marked by red dots, corresponding to points $A_{1}$ and $A_{2}$ on the spherical source.
To clearly illustrate the derivation of the amplification rate, we present the geometric relations for a radiating spherical hotspot in Fig. 9. The hotspot, denoted by $S$, has its center at $(r_{s},\,\theta_{s}=0,\,\phi_{s})$, with the initial direction of geodesics from this point indicated by the black arrow, leading to an image $I$ on the image plane. Although the example shown uses geodesics for $N=1$, the explanation applies to all image levels. The direct projection of the black hole on the image plane serves as the origin of the coordinate system, marked by a black dot, and the projection of the hotspot’s center is labeled as $P_{S}$. $\rho_{I}=\sqrt{\alpha^{2}+\beta^{2}}$ and $\rho_{S}$ represent the distances from the origin to $I$ and $P_{S}$, respectively. The gray contour line labeled CC on the image plane denotes the critical curve.
@@ V Spacetime Tomography and Hotspot Localization
The forward ray tracing method is adept at simulating images of hotspots. This section demonstrates how this method can be applied to derive observational data, specifically for measuring the BH mass, spin, inclination angle, and hotspot location based on the identification and analysis of both the direct ($N=0$) and the first lensed ($N=1$) images of a hotspot. Since the intensity of higher-order images is exponentially suppressed, our focus will primarily be on these two initial images.
Emissions near a BH can exhibit significant time variability, which may be reflected across various image levels. The intensity of these images incorporates this variability, albeit with specific time delays. Thus, by correlating intensity fluctuations across different regions on the image plane with corresponding time delays, we can confirm their common origin as emissions from the same source Hadar et al. (2021). The time delay between the two lowest image levels from a pointlike emission is sensitive to the BH mass, while their positional angle shift along the photon ring is sensitive to the BH spin Gralla and Lupsasca (2020a); Hadar et al. (2021); Andrianov et al. (2022). Additionally, the observer’s inclination angle to the BH and the hotspot’s spatial location can influence both the time delay and position angle shift, complicating the analysis as demonstrated in Sec. III.2. Forward ray tracing can mitigate these complexities by facilitating the scan over various source locations.
We model a pointlike source located at $(r_{s},\theta_{s},\phi_{s})$ outside a Kerr BH with mass $M$ and spin $a$, directing sequential geodesics toward an observer positioned at an inclination angle $\theta_{o}$ relative to the BH. Without loss of generality, we set the observer’s azimuthal angle $\phi_{o}=0$. By successfully detecting both the direct emission ($N=0$) and the first lensed image ($N=1$) of a hotspot through correlated time variability of intensity, we can ascertain precise locations of these images on the image plane, represented by
@@
where $j=a,b,c,\ldots$ labels the $j$th hotspot observed. Each selection of $\bm{\Theta}$ generates pairs of direct and the first lensed images $\bm{X}^{(j)}(\bm{\Theta})$ through the forward ray tracing method. These are then compared with the observed values $\mkern 1.0mu\overline{\mkern-1.0mu\bm{X}^{(j)}\mkern-1.0mu}\mkern 1.0mu$, which in the following two examples are assumed to match the theoretical values computed from the true parameters. Consequently, we introduce the likelihood function
@@
In scenarios where some variables in $\bm{\Theta}$ are predetermined by other measurements, such as the inclination angle $\theta_{o}$, the spin projection PA, and the SMBH mass $M$, we employ the prior probability function
@@
where $l$ labels the predetermined parameters with true values $\mkern 1.0mu\overline{\mkern-1.0mu\bm{\Theta}_{l}\mkern-1.0mu}\mkern 1.0mu$ and uncertainties $\sigma_{\bm{\Theta}_{l}}$. According to Bayes’ theorem, the posterior probability is then $p\left(\bm{\Theta}\middle|\bm{X}\right)\propto p\left(\bm{X}\middle|\bm{\Theta}\right)p(\bm{\Theta})$.
@@
V.1 M87 with Predetermined PA and $\theta_{o}$
@@ VI Discussion
A direct application of this methodology is to BH imaging, particularly in connecting a pointlike emission source outside a BH to a distant observer. This forward ray tracing approach proves more efficient than traditional backward ray tracing, especially in modeling hotspots and calculating higher-order images. By applying perturbative adjustments to the original geodesics around the emission point, our method can map a finite-size emission to distinct regions on the observer plane. Higher-order images, typically positioned near the critical curve, sequentially compress in the direction perpendicular to the curve while maintaining a stable amplification rate parallel to it. Consequently, the total area magnification rate exhibits exponential decay as the image level $N$ increases.
Our method is powerful in identifying lensed photons and conducting spacetime tomography, particularly for measuring BH mass and spin. Utilizing observations from the EHT and ngEHT that capture both direct and lensed images, our approach can precisely determine BH characteristics. For M87, with predetermined spin projection and inclination angle, a single measurement suffices to resolve the BH’s mass, spin, and the emission location. In the case of Sgr A, two independent measurements of hotspot correlations are enough to determine all relevant parameters. Given the daily occurrence of flares from hotspots around Sgr A, the prospects for successful detection are promising, especially with ngEHT, which will offer tracking of hotspot motion Emami et al. (2023) and multifrequency measurement capabilities Chael et al. (2023). The accuracy in determining universal spacetime parameters will improve as more data accumulates.
@@ Acknowledgements.
We are grateful to Xian Chen, Daniel D’Orazio, Shahar Hadar, Yosuke Mizuno, Daniel Palumbo, Diogo Ribeiro, Paul Tiede, Luka Vujeva, Bo Wang, George Wong, and Xiao Xue for useful discussions. L.Z. acknowledges financial support from Peking University for his visit to the Niels Bohr Institute. Z.Z. acknowledges financial support from China Scholarship Council (No. 202106040037). The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. V.C. and Y. C. acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. V.C. is a Villum Investigator and a DNRF Chair. V.C. acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855 and No 101131233.